# Unsourced Random Access: A Comprehensive Survey

### v2

@@ V-C MIMO Fading Channels
We now turn our focus to the URA solutions for MIMO fading channels. Several studies have investigated practical coding schemes for the MIMO URA setup, which can be categorized into five groups: CCS-based, tensor-based, pilot-based, ODMA-based, and Bayesian approaches, with inevitable overlaps between them. The MIMO URA solutions based on these approaches are reviewed in the rest of this subsection.
@@ V-C1 CCS-based schemes
In CCS-based algorithms for MIMO fading URA [97, 98], the bit sequence of each user is divided into $S$ sub-messages, which are then mapped to codewords in pilot codebooks with lower dimensions to obtain $S$ different pilots, transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot, and they are then combined to form the complete length-$B$ bit sequence for each user. To establish connections between these sub-messages, the authors in [97] incorporate parity-check bits into each sub-message, while [98] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the necessity for additional parity bits enhances the performance of [98] in comparison to [97].
@@
Eliminating the parity bits in [97] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [99, 102, 100, 101]. To stitch the decoded sub-messages, the authors of [99] employ a joint Bayesian decoding exploiting the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [100], and the similarity of the estimated channel vectors are used in [101]. The structure of the inner SPARC code allows parallel single-user outer decoding of user messages in [102], which mitigates the need for the parity check bits. An uncoupled structure is employed in [103] to design a CCS-based URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [104], where a long sub-block recovered in the first phase is used for channel estimation, and very short lengths of sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [105] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller one in both the codeword length and the number of codewords. Moreover, an uncoupled CS scheme employing polar coding and power diversity in combination with AMP-based decoding and SIC at the receiver side is proposed in [106].
V-C2 Tensor-based schemes
@@
Another trend in MIMO URA is the tensor-based schemes [107, 108, 109], where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor). For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [107], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit-sequence is detected. By employing OFDM transmission, this scheme is extended to the case with timing offsets in [108]. On the other hand, the authors of [109] propose a block-term decomposition for the decomposition of 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. The overall coding structure of tensor-based methods is depicted in Fig. 14.
@@
V-C3 Pilot-based schemes
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems, where, the idea is to divide user bit sequences into pilot and data parts. In the existing solutions, the transmitted signal of each user also consists of two parts, as depicted in Fig. 15: the pilot part which is obtained by mapping the pilot bits to a codeword in a pilot codebook (a non-orthogonal codebook [111, 12, 113, 112, 14, 110], or a codebook consisting of multiple stages of orthogonal pilots [13]), and the data part which is encoded by a channel code (a polar [13, 111, 12, 14] or an LDPC code [113, 112]). For decoding, as shown in Fig. 17, these schemes iteratively employ the following steps:
  • At the initial stage, the receiver detects the active pilots using a user activity detection technique, such as generalized OMP (gOMP) [14, 110], AMP [112, 114], multiple-measurement vector AMP (MMV-AMP) [111], energy detector [12], NP hypothesis testing [13] or CB-MLD [113].
@@ V-C3 Pilot-based schemes
  • Using the estimated channel coefficients, soft estimates of the transmitted codewords are obtained via decoupling techniques, such as MRC [111, 13, 14, 110], MMSE [12] or AMP [113], which are then used for generating log-likelihood ratio (LLR) values to be fed to single-user decoders [13, 111, 12, 113, 14, 114] or a joint factor graph [112].
  • In most pilot-based schemes, following the SIC approach, the contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the next iterations. Note that in [13, 12, 14, 110], the channel vectors are re-estimated using the successfully decoded codewords to have a better interference cancellation at the SIC stage.
@@
V-C4 ODMA-based schemes
@@
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is extended to the fading scenario as well for the case of a single-antenna receiver in [78], and a massive MIMO receiver in [110, 115, 116]. The main idea in [110] is to allocate a part of the time frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the time frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 16. At the receiver side, the gOMP algorithm is employed to detect the active pilots, followed by LMMSE channel estimation, MRC symbol estimation, single-user polar decoding, and SIC to recover the data part. Note that if the channel is a GMAC, the entire time frame is used for data transmission. On the other hand, the authors of [115] consider superimposed pilot and data transmission, where the pilot sequence occupies the entire time frame and data signal is transmitted by ODMA. The receiver operation starts with an energy detection-aided non-Bayesian pilot detection, and the same steps as in [110] are utilized for the decoding of the data part. A slotted ODMA-based scheme combining preamble transmission and convolutional LDPC coding is proposed in [117], with AMP decoding for the preamble and turbo probabilistic data association for multiuser detection of the data part at the receiver side. Furthermore, two slotted pilot-free ODMA schemes are proposed in [116] and [118]. While an iterative TIN strategy is employed for joint pattern and data detection in [116], random dictionary learning is studied in [118] in conjunction with clustering-aided activity detection.
V-C5 Bayesian approach
@@
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently. Hence, the error from one block is propagated to the next one. To resolve this issue, there is also a new set of MIMO URA schemes that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal, and their corresponding channel coefficients [120, 119]. Specifically, in [119], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Besides, the authors of [120] propose to encode the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
@@
Summary and Performance Comparison:
A summary of MIMO URA schemes is provided in Table V, accompanied by the decoding steps of the main approaches summarized in Fig. 18. Moreover, in Fig. 19, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients. The results in Fig. 19 show that due to adopting strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes in [111, 12, 13, 14, 110] usually outperform the tensor-based scheme in [107] for which a sub-optimal coding scheme is utilized, while the SKP coding-based scheme in [119], following the Bayesian approach, offers similar performance to the best performing pilot-based schemes. Finally, note that, in order to benefit from increased spatial diversity, a receiver with a massive number of antennas (i.e., a massive MIMO setup) is considered in the provided numerical examples in most MIMO URA papers. However, in order for the suitable URA candidate for next-generation applications to cover a wider range of applications, a comprehensive comparison of the state-of-the-art schemes is required, especially for a smaller number of receiver antennas.
Computational Complexity:
@@
Paper Approach Description
Gao et al., 2023 [86] - Provides achievability and converse bounds.
Fengler et al., 2021 [97] CCS-based The pilots at each slot are found using a newly proposed coordinate-wise activity detector. Then, the detected pilots of different slots are stitched together using parity bits.
Shyianov et al., 2021 [98] CCS-based Instead of employing parity-check bits (as in [97]), the correlation between different slots is used for stitching pilots, hence, the energy efficiency is improved.
@@
Jiang et al., 2021 [104] CCS-based A two-stage scheme with a long sub-message in the first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage.
@@
Decurninge et al., 2020 [107] tensor-based Users’ signal separation and detection is done by tensor decomposition.
@@
Luan et al., 2022 [109] tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [107].
Fengler et al., 2022 [111] pilot-based A pilot and a polar codeword are transmitted sequentially. For decoding, the active pilots and corresponding channel coefficients are detected using MMV-AMP and MMSE, respectively, and a polar decoder is employed for data detection.
Ahmadi et al., 2024 [13] pilot-based Each user transmits its multiple stages of orthogonal pilots along with a polar codeword through a randomly selected slot. An energy detector, an MRC, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively, followed by an LS-based SIC.
Gkagkos et al., 2023 [12] pilot-based A pilot and a random spread polar codeword are transmitted sequentially. An energy detector, an MMSE, and a polar list decoder are employed for pilot detection, channel estimation, and decoding, respectively. Finally, a SIC removes the contribution of successfully decoded signals.
Su et al., 2023 [112] pilot-based A pilot is transmitted along with an LDPC coded and modulated codeword. At the receiver side, AMP is used for pilot detection and channel estimation, and an LDPC decoder is used for data detection. The selected pilot specifies the indexes over which every symbol of the LDPC codeword is transmitted.
Su et al., 2023 [113] pilot-based A pilot and an LDPC codeword are spread and transmitted consecutively. At the receiver’s end, a maximum likelihood detection (MLD) algorithm detects active pilots, an AMP technique estimates the corresponding channels, and a SIC block is applied following the LDPC decoder.
Ozates et al., 2024 [14] pilot-based The transmission time frame is slotted, and each user randomly selects a slot to transmit a pilot along with a polar codeword. For decoding, a generalized OMP and an MMSE estimator are used for pilot detection and channel estimation, the polar decoder is used for data detection, followed by SIC.
Ozates et al., 2024 [110] pilot-based The transmission time frame is divided into pilot and data parts, and each user transmits a non-orthogonal pilot followed by its polar codeword based on ODMA. At the receiver, an iterative decoder employing gOMP, LMMSE, MRC, single-user decoding, and SIC is utilized.
Han et al., 2023 [119] Bayesian Employs FEC code, and sparse spreading.
Jiang et al., 2023 [120] Bayesian Combines sparse regression codes for encoding and a message-passing algorithm for decoding.
@@
When comparing the computational complexity of the coding strategies for MIMO URA, CCS exhibits the lowest complexity among them, as its complexity is linear in $M_{r}$ and the number of common codebook elements [97]. Tensor-based [107] and non-slotted pilot-based [12] schemes have a similar complexity that is cubic in $K_{a}$. This complexity results from the canonical polyadic decomposition in the former and MMSE filtering for channel estimation in the latter. On the other hand, the complexity of slotted pilot-based schemes [13], [14] is cubic in the number of active users in the slot, which demonstrates the complexity advantage of slotting. Furthermore, the complexity of the Bayesian approach in [119] is linear in $K_{a}$, $M_{r}$, and $n$. To summarize, for practical system parameters, CCS has the lowest complexity, while the complexities of slotted pilot-based schemes and the Bayesian approach are similar and smaller than those of tensor-based and non-slotted pilot-based schemes.
V-D Existing URA Solutions on Alternative Setups

### v3

@@ V-D MIMO Fading Channels
We now turn our focus to the URA solutions developed for MIMO fading channels, which can be categorized into five groups: CCS-based, tensor-based, pilot-based, ODMA-based, and Bayesian approaches, with inevitable overlaps among them.
@@ V-D1 CCS-based schemes
In CCS-based algorithms for URA over MIMO fading channels [105, 106], the bit sequence of each user is divided into $S$ sub-messages, which are then mapped to codewords in pilot codebooks with lower dimensions to obtain $S$ different pilots, transmitted through $S$ different slots. At the receiving end, $K_{a}$ sub-messages are detected within each slot using the ML [105] or hybrid generalized approximate message passing (HyGAMP) [106] technique, and they are then combined to form the complete length-$B$ bit sequence for each user. To establish connections among these sub-messages, the authors in [105] incorporate parity-check bits into each sub-message, while [106] relies on the correlation between channel coefficients of each user in different slots (assuming a quasi-fading channel model). Eliminating the need for additional parity bits enhances the performance of [106] compared to [105].
@@
Eliminating the parity bits in [105] turns the problem into an uncoupled compressed sensing one, which is also considered in some recent works [107, 110, 108, 109]. To stitch the decoded sub-messages, the authors in [107] employ a joint Bayesian decoding approach that exploits the channel characteristics, while geometric similarities between the codeword covariance matrices are utilized in the decoding algorithm in [108], and the similarity of the estimated channel vectors is used in [109]. The structure of the inner SPARC code allows parallel single-user outer decoding of user messages in [110], which mitigates the need for the parity check bits. An uncoupled structure is employed in [111] to design a CCS-based URA scheme for near-field communications. Furthermore, a two-phase CS coding scheme is proposed in [112], where a long sub-block recovered in the first phase is used for channel estimation, and very short-length sub-blocks are recovered in the second phase by employing a mutually orthogonal codebook and a hybrid AMP in an uncoupled structure. Moreover, the heterogeneous mobility of the massive number of users is considered in [113] via a two-codebook aided URA scheme with covariance-based maximum likelihood detection (CB-MLD) at the receiver, where the static users employ a larger codebook while the mobile users utilize a smaller one in both the codeword length and the number of codewords. Moreover, an uncoupled CS scheme employing polar coding and power diversity in combination with AMP-based decoding and SIC at the receiver side is proposed in [114].
V-D2 Tensor-based schemes
@@
Another trend in MIMO URA is the tensor-based schemes [115, 116, 117], where each user sends a rank-1 tensor of order $d$. Therefore, the received signal is a rank-$K_{a}$ tensor of order $d+1$ (considering the channel coefficient vector as an extra dimension of the tensor). For decoding, they conduct tensor decomposition on the received signal, followed by de-quantization of the tensors. In particular, in [115], by tensor decomposition, the rank-$K_{a}$ tensor of order $d+1$ is decomposed into $K_{a}$ rank-1 tensors of order $d+1$, which is equivalent to separating the signals of $K_{a}$ users, where the transmitted tensors can be identified by the first $d$ dimensions, and the $(d+1)$th dimension gives the channel vector of the user. Then, by de-quantizing the tensor of each user, its transmitted bit sequence is detected. By employing OFDM transmission, this scheme is extended to the case with timing offsets in [116]. On the other hand, the authors in [117] propose a block-term decomposition for the decomposition of 3-dimensional tensors, where two low-rank factor matrices and a factor vector are utilized in each block term. Moreover, the authors in [118] study tensor-based modulation with soft decoding, and propose a polar-coded tensor-based scheme with iterative Bayesian decoding with feedback at the receiver side. The overall coding structure of tensor-based methods is depicted in Fig. 14.
@@
V-D3 Pilot-based schemes
Pilot-based techniques constitute another category of algorithms in the context of MIMO URA systems, where the idea is to divide the user bit sequences into pilot and data parts. In the existing solutions, the transmitted signal of each user also consists of two parts, as depicted in Fig. 15: the pilot part which is obtained by mapping the pilot bits to a codeword in a pilot codebook (a non-orthogonal codebook [120, 12, 122, 121, 14, 119], or a codebook consisting of multiple stages of orthogonal pilots [13]), and the data part which is encoded by a channel code (a polar [13, 120, 12, 14] or an LDPC code [122, 121]). For decoding, as shown in Fig. 17, these schemes iteratively employ the following steps:
  • At the initial stage, the receiver detects the active pilots using a user activity detection technique, such as gOMP [14, 119], AMP [121, 124], multiple-measurement vector AMP (MMV-AMP) [120], energy detector [12], NP hypothesis testing [13], or CB-MLD [122]. Moreover, an improved version of the NNLS algorithm exploiting multiple measurements for activity detection is proposed in [123].
@@ V-D3 Pilot-based schemes
  • Using the estimated channel coefficients, soft estimates of the transmitted codewords are obtained via decoupling techniques, such as MRC [120, 13, 14, 119], MMSE [12] or AMP [122], which are then used to generate log-likelihood ratio (LLR) values to be fed to single-user decoders [13, 120, 12, 122, 14, 124] or a joint factor graph [121].
  • In most pilot-based schemes, following the SIC approach, the contribution of the successfully decoded codewords is removed from the received signal to decrease the interference level at the subsequent iterations. Note that in [13, 12, 14, 119], the channel vectors are re-estimated using the successfully decoded codewords for improved interference cancellation at the SIC stage.
@@
V-D4 ODMA-based schemes
@@
ODMA is a powerful technique that was originally proposed for URA over GMAC. However, it is also extended to the fading scenario for the case of a single-antenna receiver in [84] and a massive MIMO receiver in [119, 125, 126, 127, 129, 130, 131]. The main idea in [119] is to allocate a part of the frame for pilot transmission, which facilitates the channel estimation at the receiver side, while the rest of the frame is utilized for data transmission in an ODMA manner as illustrated in Fig. 16. At the receiver side, the gOMP algorithm is employed to detect the active pilots, followed by LMMSE channel estimation, MRC symbol estimation, single-user polar decoding, and SIC to recover the data part. Note that if the channel is a GMAC, the entire frame is used for data transmission. On the other hand, the authors in [125] consider superimposed pilot and data transmission, where the pilot sequence occupies the entire frame and the data signal is transmitted using ODMA. The receiver operation begins with energy detection-aided non-Bayesian pilot detection, and the same steps as in [119] are used for decoding the data part. A slotted ODMA-based scheme combining preamble transmission and convolutional LDPC coding is proposed in [126], with AMP decoding for the preamble and turbo probabilistic data association for multiuser detection of the data part at the receiver side. In another slotted ODMA-based scheme [127], the authors consider uncoupled ODMA where the pilot sequences and transmission patterns are selected by different bits of the message sequence. They utilize low-rank matrix factorization and simultaneous OMP [128] for activity detection and channel estimation, followed by a joint pattern and data detection based on AMP. Moreover, two slotted pilot-free ODMA schemes are proposed in [129] and [130]. While an iterative TIN strategy is employed for joint pattern and data detection in [129], random dictionary learning is studied in [130] in conjunction with clustering-aided activity detection. Finally, in a recent work [131], the authors present a strategy that is called Less is More in a slotted ODMA framework. They propose to transmit a small portion of data using a polar or LDPC code, while embedding the rest of the data into the proposed transmission structure. The transmitted signal consists of a pilot sequence followed by a polar or LDPC codeword. At the receiver side, they employ simultaneous OMP for activity detection followed by zero-forcing symbol estimation, single-user decoding, and SIC. This scheme outperforms the others for fewer than 500 active users, and its performance is slightly better than the achievability bound in [94] when the number of active users is less than 400.
V-D5 Bayesian approach
@@
In pilot-based algorithms, the channel estimation and data detection tasks are performed independently. Hence, the error from one block is propagated to the next one. To resolve this issue, a new set of MIMO URA schemes has been developed that employ a multi-layer iterative Bayesian decoder to jointly estimate each user’s transmitted signal and their corresponding channel coefficients [133, 132]. Specifically, in [132], the message bits are divided into two parts; one part is encoded by a convolutional code, and the other is encoded by index modulation. The transmitted signal is obtained by the Kronecker product of the two components. Additionally, the authors in [133] propose encoding the user payload by SPARCs, where all segments of the message bits of a user are mapped onto one codeword. This eliminates the need for an outer tree code, increasing spectral efficiency.
@@
V-E Summary and Comparison of MIMO Fading Solutions
A summary of MIMO URA schemes is provided in Table V, accompanied by the decoding steps of the main approaches summarized in Fig. 18.
V-E1 Performance Comparison
In Fig. 19, a performance comparison of the proposed coding schemes for URA with a massive MIMO receiver is provided for $n=3200$, $B=100$, $\epsilon=0.05$, and $M_{r}=50$, assuming that the channel vectors consist of i.i.d. Rayleigh fading coefficients across the different users and different antennas. The results in Fig. 19 show that, due to the adaptation of strong channel coding algorithms such as LDPC and polar codes, pilot-based schemes in [120, 12, 13, 14, 119] usually outperform the tensor-based scheme in [115] for which a relatively weak channel code is utilized, while the SKP coding-based scheme in [132], following the Bayesian approach, offers similar performance to the best-performing pilot-based solutions. Note that, to benefit from increased spatial diversity, the existing URA works on MIMO fading channels assume that the receiver is equipped with a large number of antennas (i.e., the massive MIMO regime), and only a few provide results for a small number of receive antennas [14], [132].
Taken together, the MIMO comparisons highlight that leveraging spatial dimensions is critical for scaling URA to 6G-class scenarios, which indicates that MIMO URA is among the most promising paths toward practical deployment, particularly when combined with structured coding and advanced interference management.
@@ V-E1 Performance Comparison
Paper Approach Description Synchron. vs. Asynch. Channel Estimation Channel Code
Gao et al., 2023 [94] - Provides achievability and converse bounds. - - -
Fengler et al., 2021 [105] CCS Concatenated coding scheme where the inner pilots and channel vectors are jointly recovered by ML detection and the recovered segments are stitched using a tree code. Synchron. ML -
Shyianov et al., 2021 [106] CCS CCS-based scheme where a variation of AMP is employed for inner CS decoding and channel estimation. The recovered sub-messages are stitched exploiting correlation, which improves the energy efficiency. Synchron. HyGAMP -
@@
Jiang et al., 2021 [112] CCS A two-stage scheme with a long sub-message in the first stage is used for channel estimation, and very short sub-messages are recovered using orthogonal codebooks in the second stage. Synchron. AMP -
Wang et al., 2022 [151] CCS A concatenated coding scheme employing SPARCs, tree coding and OFDM. Asynch. Embedded -
@@
Decurninge et al., 2021 [115] tensor-based Users’ signal separation and detection is done by tensor decomposition. Synchron. Embedded BCH
@@
Luan et al., 2022 [117] tensor-based A new tensor-based modulation scheme in two versions: one employing QPSK encoding, and the other implementing Grassmann encoding. It shows improved energy efficiency over the tensor-based scheme in [115]. Synchron. Embedded BCH
Fengler et al., 2022 [120] pilot-based A non-slotted scheme employing MMV-AMP pilot detection, LMMSE channel estimation, and single-user decoding. Transmit signal structure: non-orthogonal pilot + polar codeword. Synchron. LMMSE Polar
Ahmadi et al., 2024 [13] pilot-based A slotted scheme employing NP hypothesis testing pilot detection, MRC channel estimation, single-user decoding, and SIC. Transmit signal structure: Multiple stages of orthogonal pilots + polar codeword. Synchron. MRC Polar
Gkagkos et al., 2023 [12] pilot-based A non-slotted scheme employing energy test-based pilot detection, LMMSE channel estimation, single-user decoding, and SIC. Transmit signal structure: non-orthogonal pilot + randomly spread polar codeword. Synchron. LMMSE Polar
Ozates et al., 2024 [14] pilot-based A slotted scheme employing gOMP pilot detection, LMMSE channel estimation, single-user decoding, and SIC. Transmit signal structure: non-orthogonal pilot + polar codeword. Synchron. LMMSE Polar
Ozates et al., 2024 [119] ODMA A non-slotted scheme employing gOMP pilot detection, LMMSE channel estimation, single-user decoding, and SIC. Transmit signal structure: non-orthogonal pilot + polar codeword spread by ODMA. Synchron. LMMSE Polar
Han et al., 2023 [132] Bayesian SKP coding-based scheme employing BiG-AMP for outer matrix factorization and BCJR for decoding of the convolutional code. Synchron. Embedded Convol- utional
Jiang et al., 2023 [133] Bayesian A Bayesian scheme employing SPARCs for transmission and a three-layer message passing algorithm for decoding. Synchron. Embedded -
Li et al., 2025 [152] Bayesian A Bayesian scheme employing SPARCs, LDPC, and OFDM, where both timing and frequency offsets are considered. Asynch. GB-CR LDPC
V-E2 Computational Complexity Comparison
@@
Comparing the computational complexity of the coding strategies for MIMO URA, we observe that CCS exhibits the lowest complexity among them, as its complexity is linear in $M_{r}$ and the number of common codebook elements [105]. Tensor-based [115] and non-slotted pilot-based [12] schemes have a similar complexity (cubic in $K_{a}$). This complexity results from the canonical polyadic decomposition in the former and MMSE filtering for channel estimation in the latter. On the other hand, the complexity of slotted pilot-based schemes [13], [14] is cubic in the number of active users in the slot, which demonstrates the complexity advantage of slotting. Furthermore, the complexity of the Bayesian approach in [132] is linear in $K_{a}$, $M_{r}$, and $n$. To summarize, for practical system parameters, CCS has the lowest complexity, while the complexities of slotted pilot-based schemes and the Bayesian approach are similar and smaller than those of tensor-based and non-slotted pilot-based approaches.
V-F Existing URA Solutions on Alternative Setups